# Gaussian Shading: Provable Performance-Lossless Image Watermarking for Diffusion Models

### v1

@@ 3.4 Proof of Lossless Performance
where $\delta$ is non-negligible with respect to the key $K$. Let the iterative denoising process be denoted as $Q(\cdot)$, and substitute the LDM encoder $\mathcal{E}$ into Eq. 7, we have
$\left|\Pr\left[\mathcal{A}\left(\mathcal{E}\left(Q\left(z_{T}^{s}\right)\right)\right)=1|m=E\left(K,s^{d}\right)\right]\right.$ (8)
$\left.\quad-\Pr\left[\mathcal{A}\left(\mathcal{E}\left(Q\left(z_{T}\right)\right)\right)=1|z_{T}\leftarrow\mathcal{N}\left(0,I\right)\right]\right|=\delta,$
@@
$\left|\Pr[\mathcal{A}\left(\mathcal{E}\left(Q\left(S\left(m\right)\right)\right)\right)=1|m=E(K,s^{d})]\right.$ (9)
$\left.\quad\quad\quad\quad\quad\quad-\Pr\left[\mathcal{A}\left(\mathcal{E}\left(Q\left(S\left(r\right)\right)\right)\right)=1\right]\right|=\delta.$
Sampling $S(\cdot)$, denoising $Q(\cdot)$, and encoder $\mathcal{E}$ can be considered as subroutines that the tester $\mathcal{A}_{\mathcal{E},Q,S}$ can use. Thus, Eq. 9 can be simplified,
$\left|\Pr[\mathcal{A}_{\mathcal{E},Q,S}\left(m\right)=1|m=E\left(K,s^{d}\right)]\right.$ (10)
$\left.\quad\quad\quad\quad\quad\quad\quad\quad-\Pr\left[\mathcal{A}_{\mathcal{E},Q,S}\left(r\right)=1\right]\right|=\delta.$
Note that $S(\cdot)$, $Q(\cdot)$, and $\mathcal{E}$ are all polynomial-time programs, so the time taken by the tester $\mathcal{A}_{\mathcal{E},Q,S}$ to make the distinction is also polynomial. Eq. 10 essentially states that it is possible to distinguish between $m$ and $r$ in polynomial time. However, we have used the computationally secure stream cipher ChaCha20 [3] in watermark randomization, which means that $m$ as a pseudorandom sequence cannot be distinguished from a truly random sequence in polynomial time. Eq. 10 contradicts the computational security property of ChaCha20 [3]. Therefore, Eq. 10 is not valid, leading us back to our initial assumption that Eq. 7 is also not valid. This implies that the watermarked image $X^{s}$ and the normally generated image $X$ are indistinguishable in polynomial time. Hence, Gaussian Shading is performance-lossless under chosen watermark tests.
@@ 4.4 Ablation Studies
Noise intensities. To further test the robustness, we conduct experiments using different intensities of noises. See Figs. 5(b), 5(c), 5(d), 5(f), 5(g), 5(h), 5(i), 5(e) and 5(j), for Random Crop and Gaussian Noise, performance declines significantly with higher intensities. However, for the other seven types of noise, even at high intensities, the bit accuracy remains approximately $80\%$.
@@ 4.5 Attacks against Gaussian Shading
Compression attack. We utilize popular auto-encoders [2, 5, 11, 33] to compare Stable Signature (SS) with Gaussian Shading across various compression rates. Additionally, we assess the compression quality through the PSNR between the compressed and watermarked images. See Figs. 6(a) and 6(b), Gaussian Shading significantly outperforms Stable Signature. This is because Gaussian Shading diffuses the watermark across the entire semantic space of images, while Stable Signature relies solely on the image texture.
@@ 8.4 Additional Visual Results
Original
DwtDct [7]
DwtDctSvd [7]
RivaGAN [49]
Stable Signature [12]
Ours
@@
Prompt
Original
Tree-Ring [44]
Ours
A bird is sitting on a bowl of birdseed.
A man holding open an oven door in a kitchen.
A skillet on a stove with vegetables in it.
This is two birds pecking at the remnants of a burger at an outdoor restaurant.
Many surfboards are propped against a rail on the beach.
Figure 10: Visual comparison between Tree-Ring and Gaussian Shading on prompts of the validation set of MS-COCO-2017, at resolution 512. In contrast to the original model and our Gaussian Shading, Tree-Ring alters the distribution of the latent representations, potentially resulting in the generation of images characterized by semantic inconsistencies or diminished quality. This figure illustrates an instance of such a case, where the Gaussian Shading preserves t.he distribution, thereby avoiding this issue.

### v3

@@ 3.4 Proof of Lossless Performance
where $\delta$ is non-negligible with respect to the key $K$. Let the iterative denoising process be denoted as $Q(\cdot)$, and substitute the LDM decoder $\mathcal{D}$ into Eq. 7, we have
$\left|\Pr\left[\mathcal{A}\left(\mathcal{D}\left(Q\left(z_{T}^{s}\right)\right)\right)=1|m=E\left(K,s^{d}\right)\right]\right.$ (8)
$\left.\quad-\Pr\left[\mathcal{A}\left(\mathcal{D}\left(Q\left(z_{T}\right)\right)\right)=1|z_{T}\leftarrow\mathcal{N}\left(0,I\right)\right]\right|=\delta,$
@@
$\left|\Pr[\mathcal{A}\left(\mathcal{D}\left(Q\left(S\left(m\right)\right)\right)\right)=1|m=E(K,s^{d})]\right.$ (9)
$\left.\quad\quad\quad\quad\quad\quad-\Pr\left[\mathcal{A}\left(\mathcal{D}\left(Q\left(S\left(r\right)\right)\right)\right)=1\right]\right|=\delta.$
Sampling $S(\cdot)$, denoising $Q(\cdot)$, and decoder $\mathcal{D}$ can be considered as subroutines that the tester $\mathcal{A}_{\mathcal{D},Q,S}$ can use. Thus, Eq. 9 can be simplified,
$\left|\Pr[\mathcal{A}_{\mathcal{D},Q,S}\left(m\right)=1|m=E\left(K,s^{d}\right)]\right.$ (10)
$\left.\quad\quad\quad\quad\quad\quad\quad\quad-\Pr\left[\mathcal{A}_{\mathcal{D},Q,S}\left(r\right)=1\right]\right|=\delta.$
Note that $S(\cdot)$, $Q(\cdot)$, and $\mathcal{D}$ are all polynomial-time programs, so the time taken by the tester $\mathcal{A}_{\mathcal{D},Q,S}$ to make the distinction is also polynomial. Eq. 10 essentially states that it is possible to distinguish between $m$ and $r$ in polynomial time. However, we have used the computationally secure stream cipher ChaCha20 [3] in watermark randomization, which means that $m$ as a pseudorandom sequence cannot be distinguished from a truly random sequence in polynomial time. Eq. 10 contradicts the computational security property of ChaCha20 [3]. Therefore, Eq. 10 is not valid, leading us back to our initial assumption that Eq. 7 is also not valid. This implies that the watermarked image $X^{s}$ and the normally generated image $X$ are indistinguishable in polynomial time. Hence, Gaussian Shading is performance-lossless under chosen watermark tests.
@@ 4.4 Ablation Studies
Noise intensities. To further test the robustness, we conduct experiments using different intensities of noises. See Figs. 6(b), 6(c), 6(d), 6(f), 6(g), 6(h), 6(i), 6(e) and 6(j), for Random Crop and Gaussian Noise, performance declines significantly with higher intensities. However, for the other seven types of noise, even at high intensities, the bit accuracy remains approximately $80\%$.
@@ 4.5 Attacks against Gaussian Shading
Compression attack. We utilize popular auto-encoders [2, 5, 11, 33] to compare Stable Signature (SS) with Gaussian Shading across various compression rates. Additionally, we assess the compression quality through the PSNR between the compressed and watermarked images. See Figs. 7(a) and 7(b), Gaussian Shading significantly outperforms Stable Signature. This is because Gaussian Shading diffuses the watermark across the entire semantic space of images, while Stable Signature relies solely on the image texture.
@@ 8.4 Additional Visual Results
Original Refer to caption Refer to caption
DwtDct [7] Refer to caption Refer to caption Refer to caption Refer to caption
DwtDctSvd [7] Refer to caption Refer to caption Refer to caption Refer to caption
RivaGAN [49] Refer to caption Refer to caption Refer to caption Refer to caption
Stable Signature [12] Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption
@@
Prompt Original Tree-Ring [44] Ours
A bird is sitting on a bowl of birdseed. Refer to caption Refer to caption Refer to caption
A man holding open an oven door in a kitchen. Refer to caption Refer to caption Refer to caption
A skillet on a stove with vegetables in it. Refer to caption Refer to caption Refer to caption
This is two birds pecking at the remnants of a burger at an outdoor restaurant. Refer to caption Refer to caption Refer to caption
Many surfboards are propped against a rail on the beach. Refer to caption Refer to caption Refer to caption
Figure 10: Visual comparison between Tree-Ring and Gaussian Shading on prompts of the validation set of MS-COCO-2017, at resolution 512. Two methods are applied with the same input latent representations. In contrast to the original model and our Gaussian Shading, Tree-Ring alters the distribution of the latent representations, potentially resulting in the generation of images characterized by semantic inconsistencies or diminished quality. This figure illustrates an instance of such a case, where the Gaussian Shading preserves the distribution, thereby avoiding this issue.